OpenStudy (chrisneedshelp):

What is the equation of the line, in slope-intercept form, that passes through (4, 2) and (-2, -3)? 5x - 6y - 8 = 0 3y = -2x + 20 y = x - 5 over 6 times - 4 over 3

OpenStudy (anonymous):

equation as in the form of y=mx+c or ax+y+c=0?

OpenStudy (anonymous):

First you get the slope of the equation; \(m = \dfrac {y-y_1}{x-x_1}\), then you use rewrite the whole thing in slope-intercept form \(y = mx +b\)

OpenStudy (anonymous):

ecah1, you'll have something like this \(y + 3 = \dfrac{5}{6}(x+2) \longrightarrow y = \frac{5}{6}x + \frac{5}{3}-3 = \frac{5}{6}x - \frac{4}{3}\)

OpenStudy (anonymous):

sure your right dexter.. my bad.. sowi :( i see wat i forgot to put include in my calculations. and yes your ryt :D